OpenStudy (anonymous):

How do you evaluate the rational exponent of 121 1/2?

OpenStudy (anonymous):

wait...did you mean 121^1/2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that would be 11.

OpenStudy (anonymous):

You have to think to yourself, "What two numbers can I multiply to get to 121?" the 1/2 power means find two numbers that multiply to get 121

OpenStudy (anonymous):

Find a number that, if you multiply it by itself, you get 121

OpenStudy (anonymous):

11 x 11 = 121

OpenStudy (anonymous):

\[\sqrt{121}=11\]

OpenStudy (anonymous):

121 1/2

OpenStudy (anonymous):

Would that be considered a "real number"?

OpenStudy (anonymous):

Yes, 11 is a real number.

OpenStudy (anonymous):

what do we do with the half when showing our work?

OpenStudy (anonymous):

Using the calculator it comes up with 11.02270384

OpenStudy (anonymous):

\[x^{1/2}=\sqrt{x}\]

OpenStudy (anonymous):

Some calculator has error.

OpenStudy (anonymous):

Google calc says the same answer

OpenStudy (anonymous):

Can you tell me the website of Google calc?

OpenStudy (anonymous):

Did you enter it into the calculator correctly?

OpenStudy (anonymous):

\[\sqrt{121.5}=11.02270384\]

OpenStudy (anonymous):

\[\sqrt{121}=11\]

OpenStudy (anonymous):

I typed in 121 pressed the a b/c button then 1 followed by a b/c then 2 then pressed the square root button. This resulted in the above number.

OpenStudy (anonymous):

I think you should type 121, then press square root button, and get 11.

OpenStudy (anonymous):

121 1/2 is written as a fraction and not as an exponent. Would that make a diffence?

OpenStudy (anonymous):

oh, if you mean (121 1/2)^(1/2), the answer is 11.02

OpenStudy (anonymous):

\[121\frac{1}{2}\]?

OpenStudy (anonymous):

you want rewrite in fraction?